# III. Covariant canonical quantum gravity models

### v2

@@ I Introduction
Nearly all earlier studies on PBHs as DM candidates worked under the assumption of these objects being Schwarzschild or Kerr PBHs. Under such an assumption, it has been found that PBHs can account for the entire DM component of the Universe only within the so-called “asteroid mass window”, for PBH masses $10^{17}\,{\text{g}}\lesssim M_{\text{pbh}}\lesssim 10^{23}\,{\text{g}}$: lighter PBHs are tightly constrained by the non-observation of products of their Hawking evaporation, whereas heavier ones are strongly constrained by the absence of microlensing signatures expected from their presence in the galactic halo Bai and Orlofsky (2019); Smyth et al. (2020); Coogan et al. (2021); Ray et al. (2021); Auffinger (2022); Ghosh and Mishra (2024); Miller et al. (2022); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2024); Gorton and Green (2024); Dent et al. (2025); Tamta et al. (2025); Tinyakov (2024); Loeb (2024); Li et al. (2025a). All observational constraints on PBHs, including those determining the extension of the asteroid mass window, are contingent on the assumption of Schwarzschild and/or Kerr PBHs. While on the phenomenological side this is a reasonable assumption, on the theoretical side the situation is quite different, given that both space-times feature well-known curvature singularities. These considerations motivated some of us, in two earlier companion papers Calzà et al. (2025a, b), to entertain the possibility that DM is composed of primordial regular (non-singular) BHs (PRBHs):  [^1] Specifically, in the first work Calzà et al. (2025a) we considered three tr (time-radial)-symmetric regular space-times (the Bardeen Bardeen (1968), Hayward Hayward (2006), and Culetu-Ghosh-Simpson-Visser ones Culetu (2013, 2015); Ghosh (2015); Simpson and Visser (2019a)), whereas the second Calzà et al. (2025b) studied three non-tr-symmetric metrics (the Simpson-Visser Simpson and Visser (2019b), Peltola-Kunstatter Peltola and Kunstatter (2009a, b), and D’Ambrosio-Rovelli ones Bianchi et al. (2018); D’Ambrosio and Rovelli (2018)). See e.g. Refs. Easson (2003); Dymnikova and Khlopov (2015); Pacheco (2018); Arbey and Auffinger (2021); Arbey et al. (2022); de Freitas Pacheco et al. (2023); Banerjee et al. (2025); Davies et al. (2025); Dialektopoulos et al. (2025); Trivedi and Loeb (2025); Carr et al. (2025); Aşmanoğlu et al. (2025); Jusufi and Singleton (2025) for other works considering the possibility that primordial non-Schwarzschild/Kerr objects, and/or remnants thereof, may play a role in the DM problem. in short, we found that evaporation constraints on all the six PRBH space-times studied are significantly relaxed compared to their Schwarzschild counterparts (and can actually be completely evaded under certain extremality assumptions), resulting in a much larger window in mass where PRBHs can account for all of the DM compared to the Schwarzschild PBHs case.
@@ II Zhang-Lewandowski-Ma-Yang regular black hole
The Penrose-Hawking singularity theorems have established the unavoidable appearance of singularities in GR Penrose (1965); Hawking and Penrose (1970) (see Refs. Ong (2020); de Haro et al. (2023); Trivedi (2024) for recent reviews). These singularities are clearly undesirable, since they signal a breakdown of predictability (see, however, Refs. Sachs et al. (2021); Ashtekar et al. (2021, 2022) for a different view on this problem): for this reason, significant effort has been devoted to the study of so-called regular BHs, i.e. space-times which are free of singularities Borde (1997); Ayon-Beato and Garcia (1998, 1999); Bronnikov and Fabris (2006); Berej et al. (2006); Bronnikov et al. (2012); Rinaldi (2012); Stuchlík and Schee (2014); Schee and Stuchlik (2015); Johannsen (2013); Myrzakulov et al. (2016); Fan and Wang (2016); Sebastiani et al. (2017); Toshmatov et al. (2017); Chinaglia and Zerbini (2017); Frolov (2018); Bertipagani et al. (2021); Nashed and Saridakis (2022); Simpson and Visser (2022); Franzin et al. (2022); Chataignier et al. (2023); Ghosh et al. (2023); Khodadi and Pourkhodabakhshi (2022); Farrah et al. (2023); Fontana and Rinaldi (2023); Boshkayev et al. (2023); Luongo et al. (2023); Luongo and Quevedo (2024); Cadoni et al. (2023); Giambò and Luongo (2024); Cadoni et al. (2024); Luongo et al. (2024); Sajadi et al. (2024); Javed et al. (2024); Ditta et al. (2024); Al-Badawi et al. (2024); Övgün et al. (2024); Corona et al. (2024); Bueno et al. (2025a); Konoplya and Zhidenko (2024); Pedrotti and Vagnozzi (2024); Bronnikov (2024); Kurmanov et al. (2024); Bolokhov et al. (2024); Agrawal et al. (2024); Belfiglio et al. (2025); Stashko (2024); Faraoni and Rinaldi (2024); Khodadi and Firouzjaee (2024); Calzà et al. (2024); Estrada and Aros (2025); Kumar Walia (2024); Li et al. (2025b); Estrada (2025); Benavides-Gallego et al. (2024); Frolov et al. (2025); Balart et al. (2025); Bueno et al. (2025b, c); Vertogradov et al. (2025); Sajadi et al. (2025); Xiong et al. (2025); Estrada et al. (2025); Casadio et al. (2025); Harada et al. (2025); Bueno et al. (2025d); Fauzi et al. (2025); Kala et al. (2025); Capozziello et al. (2025); Pedrotti and Calzà (2025); Urmanov et al. (2025); Pinto et al. (2025); Calzá et al. (2025); Bueno et al. (2025e); Neves (2025); Eichhorn and Fernandes (2025); Bonanno et al. (2025); Bueno et al. (2025f), see Refs. Sebastiani and Zerbini (2022); Lan et al. (2023); Carballo-Rubio et al. (2025) for reviews.  [^2] It is worth clarifying that finiteness of curvature invariants is not in itself a guarantee of geodesic completeness. This is a problem which affects a number of well-known regular BH solutions Zhou and Modesto (2023). In addition, an issue currently under debate concerns whether several regular BH solutions are unstable against perturbations, due to the so-called mass inflation instability Carballo-Rubio et al. (2021, 2022); Bonanno et al. (2023); Bonanno and Saueressig (2022); Carballo-Rubio et al. (2023). In this work, following up on our earlier companion papers Calzà et al. (2025a, b), we will consider the possibility of primordial regular BHs making up the DM. The generic space-time we consider is characterized by a line element taking the following form:
@@ IV Results and discussion
Four comments are in order at this point. Firstly, we note that the results reported go completely in the opposite direction compared to our two earlier works Calzà et al. (2025a, b), where we observed weaker constraints on $f_{\text{pbh}}$ and thereby a broadening of the asteroid mass window. Although this was probably not sufficiently emphasized in Refs. Calzà et al. (2025a, b), the evaporation constraints for the metrics considered there can in fact be completely eliminated if the regularizing parameter reaches its extremal value, in which case the regular BH temperature reaches zero, therefore leading to a non-evaporating relic (see also Ref. Davies et al. (2025)). It is nevertheless worth reminding the reader that the metrics considered earlier were all phenomenological or, at best, quantum gravity-inspired. The metric we consider here is instead an exact solution of a generally covariant effective canonical (loop) quantum gravity model, which is not only non-singular but also free of Cauchy horizons, making it a serious candidate to address the singularity problem in a theoretically healthy way. The constraints on $f_{\text{pbh}}$ derived here are therefore arguably somewhat more robust and theoretically well-grounded compared to the earlier ones.
In addition, as in our earlier works Calzà et al. (2025a, b), we are considering spherically symmetric space-times, i.e. non-rotating BHs. In the Schwarzschild case, considering the PBH spin and therefore the Kerr metric has been demonstrated to lead to tighter constraints on $f_{\text{pbh}}$, thereby leading to a smaller asteroid mass window Arbey et al. (2020, 2021b). The main reason is that, even though increasing the Kerr BH spin lowers its temperature, the GBF associated to the maximally co-rotating $m$ mode starts increasing at much smaller energies compared to the other modes, leading to an overall enhancement of the Hawking evaporation spectrum. By thermodynamical considerations, it is more than plausible that including rotation in the ZLMY BH would lead to a similar effect, and therefore that the asteroid mass window would be closed down even further compared to what is shown in Fig. 4. However, checking this explicitly would require one to construct the rotating version of Eq. (3), a task to which the Newman-Janis algorithm is not well-suited given the non-phenomenological nature of the metric. We therefore defer this interesting question to follow-up work.  [^4] In a similar spirit, it might be interesting to assess whether potential new windows for small $M_{\text{pbh}}\lesssim 10^{15}\,{\text{g}}$ PBHs as DM candidates opened up by the memory burden effect, which suppresses the evaporation of BHs once about half of their mass has been lost Dvali et al. (2020), may be altered (relative to the Schwarzschild case) when working with primordial regular BHs.
@@
Finally, we remind the reader of a number of caveats surrounding our results, already discussed in Refs. Calzà et al. (2025a, b). Firstly, we are treating the parameter $\xi/M$ as an “universal hair” (in the language of Ref. Vagnozzi et al. (2023)), but whether or not this is motivated from a fundamental perspective is an open question which should eventually be addressed. In addition, our statements on the asteroid mass window shrinking are contingent on the upper edge thereof, which is determined by lensing constraints, remaining fixed. As argued earlier Calzà et al. (2025a, b), we do indeed expect this to be the case, although a detailed study of microlensing constraints on primordial regular BHs (including primordial ZLMY BHs) will be the subject of a separate work. Finally, non-evaporation constraints on $f_{\text{pbh}}$ (e.g. dynamical, accretion, and CMB limits, see Ref. Carr et al. (2024) for a recent summary) are expected to be relevant within a completely different mass range, and will be the subject of a follow-up work.
@@ V Conclusions
Together with our earlier works, our results further support the idea that there is a promising common ground between the DM, singularity, and QG problems, worthy of further exploration. As remarked in Refs. Calzà et al. (2025a, b), this series of works should be intended as pilot studies in this exciting direction, as several open problems remain. For instance, an important open question is whether the formation mechanism of primordial regular BHs is significantly different from the Schwarzschild case, as departures from the assumptions underlying Birkhoff’s theorem may entail a non-unique endpoint for gravitational collapse: this may potentially lead to complementary observational signatures, which may offer new routes towards testing the regular nature of astrophysical BHs. Moreover, an important task is to revisit non-evaporation (lensing, accretion, dynamical) constraints on primordial regular BHs: although we have loosely argued that we do not expect these to change significantly, such a statement remains to be checked explicitly. While we defer all these interesting points to follow-up work, we stress once more a general but important message of the present work: even ideas or frameworks which at first may appear remote or speculative – such as LQG – can have important consequences for both astrophysical observations and other open problems in physics, whereas the issue of a theoretically consistent formulation can make a significant difference at the level of phenomenological predictions.

### v3

@@ I Introduction
Nearly all earlier studies on PBHs as DM candidates worked under the assumption of these objects being Schwarzschild or Kerr PBHs. Under such an assumption, it has been found that PBHs can account for the entire DM component of the Universe only within the so-called “asteroid mass window”, for PBH masses $10^{17}\,{\text{g}}\lesssim M_{\text{pbh}}\lesssim 10^{23}\,{\text{g}}$: lighter PBHs are tightly constrained by the non-observation of products of their Hawking evaporation, whereas heavier ones are strongly constrained by the absence of microlensing signatures expected from their presence in the galactic halo Bai and Orlofsky (2019); Smyth et al. (2020); Coogan et al. (2021); Ray et al. (2021); Auffinger (2022); Ghosh and Mishra (2024); Miller et al. (2022); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2024); Gorton and Green (2024); Dent et al. (2025); Tamta et al. (2025); Tinyakov (2024); Loeb (2024); Li et al. (2025a). All observational constraints on PBHs, including those determining the extension of the asteroid mass window, are contingent on the assumption of Schwarzschild and/or Kerr PBHs. While on the phenomenological side this is a reasonable assumption, on the theoretical side the situation is quite different, given that both space-times feature well-known curvature singularities. These considerations motivated some of us, in two earlier companion papers Calzà et al. (2025a, b), to entertain the possibility that DM is composed of primordial regular (non-singular) BHs (PRBHs):  [^1] Specifically, in the first work Calzà et al. (2025a) we considered three tr (time-radial)-symmetric regular space-times (the Bardeen Bardeen (1968), Hayward Hayward (2006), and Culetu-Ghosh-Simpson-Visser ones Culetu (2013, 2015); Ghosh (2015); Simpson and Visser (2019a)), whereas the second Calzà et al. (2025b) studied three non-tr-symmetric metrics (the Simpson-Visser Simpson and Visser (2019b), Peltola-Kunstatter Peltola and Kunstatter (2009a, b), and D’Ambrosio-Rovelli ones Bianchi et al. (2018); D’Ambrosio and Rovelli (2018)). See e.g. Refs. Easson (2003); Dymnikova and Khlopov (2015); Pacheco (2018); Arbey and Auffinger (2021); Arbey et al. (2022); de Freitas Pacheco et al. (2023); Banerjee et al. (2025); Davies et al. (2025); Dialektopoulos et al. (2025); Trivedi and Loeb (2025a); Carr et al. (2025); Aşmanoğlu et al. (2025); Jusufi and Singleton (2025); Khodadi (2025); Trivedi and Loeb (2025b); Loc (2025) for other works considering the possibility that primordial non-Schwarzschild/Kerr objects, and/or remnants thereof, may play a role in the DM problem. in short, we found that evaporation constraints on all the six PRBH space-times studied are significantly relaxed compared to their Schwarzschild counterparts (and can actually be completely evaded under certain extremality assumptions), resulting in a much larger window in mass where PRBHs can account for all of the DM compared to the Schwarzschild PBHs case.
@@ II Zhang-Lewandowski-Ma-Yang regular black hole
The Penrose-Hawking singularity theorems have established the unavoidable appearance of singularities in GR Penrose (1965); Hawking and Penrose (1970) (see Refs. Ong (2020); de Haro et al. (2023); Trivedi (2024) for recent reviews). These singularities are clearly undesirable, since they signal a breakdown of predictability (see, however, Refs. Sachs et al. (2021); Ashtekar et al. (2021, 2022) for a different view on this problem): for this reason, significant effort has been devoted to the study of so-called regular BHs, i.e. space-times which are free of singularities Borde (1997); Ayon-Beato and Garcia (1998, 1999); Bronnikov and Fabris (2006); Berej et al. (2006); Bronnikov et al. (2012); Rinaldi (2012); Stuchlík and Schee (2014); Schee and Stuchlik (2015); Johannsen (2013); Myrzakulov et al. (2016); Fan and Wang (2016); Sebastiani et al. (2017); Toshmatov et al. (2017); Chinaglia and Zerbini (2017); Frolov (2018); Bertipagani et al. (2021); Nashed and Saridakis (2022); Simpson and Visser (2022); Franzin et al. (2022); Chataignier et al. (2023); Ghosh et al. (2023); Khodadi and Pourkhodabakhshi (2022); Farrah et al. (2023); Fontana and Rinaldi (2023); Boshkayev et al. (2023); Luongo et al. (2023); Luongo and Quevedo (2024); Cadoni et al. (2023); Giambò and Luongo (2024); Cadoni et al. (2024); Luongo et al. (2024); Sajadi et al. (2024); Javed et al. (2024); Ditta et al. (2024); Al-Badawi et al. (2024); Övgün et al. (2024); Corona et al. (2024); Bueno et al. (2025a); Konoplya and Zhidenko (2024); Pedrotti and Vagnozzi (2024); Bronnikov (2024); Kurmanov et al. (2024); Capozziello et al. (2024); Bolokhov et al. (2024); Agrawal et al. (2024); Belfiglio et al. (2025); Stashko (2024); Faraoni and Rinaldi (2024); Khodadi and Firouzjaee (2024); Calzà et al. (2024); Estrada and Aros (2025); Kumar Walia (2024); Li et al. (2025b); Estrada (2025); Benavides-Gallego et al. (2024); Frolov et al. (2025); Balart et al. (2025); Bueno et al. (2025b, c); Vertogradov et al. (2025); Sajadi et al. (2025); Xiong et al. (2025); Estrada et al. (2025); Casadio et al. (2025); Harada et al. (2025); Bueno et al. (2025d); Fauzi et al. (2025); Kala et al. (2025); Capozziello et al. (2025a); Pedrotti and Calzà (2025); Urmanov et al. (2025); Pinto et al. (2025); Calzá et al. (2025); Bueno et al. (2025e); Neves (2025); Capozziello et al. (2025b); Eichhorn and Fernandes (2025); Bonanno et al. (2025); Bueno et al. (2025f); Calzà et al. (2025), see Refs. Sebastiani and Zerbini (2022); Lan et al. (2023); Carballo-Rubio et al. (2025) for reviews.  [^2] It is worth clarifying that finiteness of curvature invariants is not in itself a guarantee of geodesic completeness. This is a problem which affects a number of well-known regular BH solutions Zhou and Modesto (2023). In addition, an issue currently under debate concerns whether several regular BH solutions are unstable against perturbations, due to the so-called mass inflation instability Carballo-Rubio et al. (2021, 2022); Bonanno et al. (2023); Bonanno and Saueressig (2022); Carballo-Rubio et al. (2023). In this work, following up on our earlier companion papers Calzà et al. (2025a, b), we will consider the possibility of primordial regular BHs making up the DM. The generic space-time we consider is characterized by a line element taking the following form:
@@ IV Results and discussion
Five comments are in order at this point. Firstly, we note that the results reported go completely in the opposite direction compared to our two earlier works Calzà et al. (2025a, b), where we observed weaker constraints on $f_{\text{pbh}}$ and thereby a broadening of the asteroid mass window. Although this was probably not sufficiently emphasized in Refs. Calzà et al. (2025a, b), the evaporation constraints for the metrics considered there can in fact be completely eliminated if the regularizing parameter reaches its extremal value, in which case the regular BH temperature reaches zero, therefore leading to a non-evaporating relic (see also Ref. Davies et al. (2025)). It is nevertheless worth reminding the reader that the metrics considered earlier were all phenomenological or, at best, quantum gravity-inspired. The metric we consider here is instead an exact solution of a generally covariant effective canonical (loop) quantum gravity model, which is not only non-singular but also free of Cauchy horizons, making it a serious candidate to address the singularity problem in a theoretically healthy way. The constraints on $f_{\text{pbh}}$ derived here are therefore arguably somewhat more robust and theoretically well-grounded compared to the earlier ones.
Next, as in our earlier works Calzà et al. (2025a, b), we are considering spherically symmetric space-times, i.e. non-rotating BHs. In the Schwarzschild case, considering the PBH spin and therefore the Kerr metric has been demonstrated to lead to tighter constraints on $f_{\text{pbh}}$, thereby leading to a smaller asteroid mass window Arbey et al. (2020, 2021b). The main reason is that, even though increasing the Kerr BH spin lowers its temperature, the GBF associated to the maximally co-rotating $m$ mode starts increasing at much smaller energies compared to the other modes, leading to an overall enhancement of the Hawking evaporation spectrum. By thermodynamical considerations, it is more than plausible that including rotation in the ZLMY BH would lead to a similar effect, and therefore that the asteroid mass window would be closed down even further compared to what is shown in Fig. 4. However, checking this explicitly would require one to construct the rotating version of Eq. (3), a task to which the Newman-Janis algorithm is not well-suited given the non-phenomenological nature of the metric. We therefore defer this interesting question to follow-up work.  [^4] In a similar spirit, it might be interesting to assess whether potential new windows for small $M_{\text{pbh}}\lesssim 10^{15}\,{\text{g}}$ PBHs as DM candidates opened up by the memory burden effect, which suppresses the evaporation of BHs once about half of their mass has been lost Dvali et al. (2020), may be altered (relative to the Schwarzschild case) when working with primordial regular BHs.
@@
Additionally, we remind the reader of a number of caveats surrounding our results, already discussed in Refs. Calzà et al. (2025a, b). Firstly, we are treating the parameter $\xi/M$ as an “universal hair” (in the language of Ref. Vagnozzi et al. (2023)), but whether or not this is motivated from a fundamental perspective is an open question which should eventually be addressed. In addition, our statements on the asteroid mass window shrinking are contingent on the upper edge thereof, which is determined by lensing constraints, remaining fixed. As argued earlier Calzà et al. (2025a, b), we do indeed expect this to be the case, although a detailed study of microlensing constraints on primordial regular BHs (including primordial ZLMY BHs) will be the subject of a separate work. Finally, non-evaporation constraints on $f_{\text{pbh}}$ (e.g. dynamical, accretion, and CMB limits, see Ref. Carr et al. (2024) for a recent summary) are expected to be relevant within a completely different mass range, and will be the subject of a follow-up work.
Beyond these phenomenological aspects, it is worth briefly commenting on the broader and more general cosmological and astrophysical implications of PBH DM. On large scales, a population of PBHs (singular or regular) behaves dynamically as a pressureless, collisionless medium. The discrete nature of the PBH distribution, while potentially leaving interesting imprints, becomes irrelevant once averaged over many objects: therefore, prior to horizon re-entry, PBHs generate gravitational potentials which set the stage for the later growth of baryon perturbations, in complete analogy to standard cold DM. In this respect, we note that the cosmological evolution driven by primordial regular BHs should be practically indistinguishable from that produced by their singular counterparts.
The discrete nature of the PBH distribution may nevertheless lead to interesting astrophysical signatures. Such a discreteness is inevitably accompanied by Poisson fluctuations, which introduce additional (shot noise-like) small-scale power. This would influence the angular momentum evolution of early halos as well as their collapse dynamics and could, at least qualitatively, be related to the long-standing galactic spin problem. In addition, the clustering of PBHs on galactic scales could plausibly affect the observed morphology-density relation of galaxies. For instance, denser PBH environments might favor merger-driven, pressure-supported systems such as elliptical and lenticular galaxies, with the converse true in less dense environments, which may aid the survival of rotationally supported disks, thereby leading to a preferential formation of (flocculent) spiral and irregular galaxies. While these deliberately qualitative cosmological and astrophysical considerations lie beyond the scope of this work, they show how the phenomenology of primordial (singular or regular) BHs, including ZLMY PBHs, fit into the broader cosmological and astrophysical picture.
@@ V Conclusions
Together with our earlier works, our results further support the idea that there is a promising common ground between the DM, singularity, and QG problems, worthy of further exploration. As remarked in Refs. Calzà et al. (2025a, b), this series of works should be intended as pilot studies in this exciting direction, as several open problems remain. For instance, an important open question is whether the formation mechanism of primordial regular BHs is significantly different from the Schwarzschild case, as departures from the assumptions underlying Birkhoff’s theorem may entail a non-unique endpoint for gravitational collapse: this may potentially lead to complementary observational signatures, which may offer new routes towards testing the regular nature of astrophysical BHs. Another key task is to revisit non-evaporation (lensing, accretion, dynamical) constraints on primordial regular BHs: although we have loosely argued that we do not expect these to change significantly, such a statement remains to be checked explicitly. Finally, we note that the same canonical polymerized framework underlying the ZLMY space-time, which bounds curvature invariants and cures future spacelike singularities, might also tame future timelike ones (of more cosmological interest) de Haro et al. (2023): indeed, works on loop quantum cosmology have shown how future finite-time singularities (e.g. Big Rip/Type I, as well as often Type III) are generally avoided in these scenarios, whereas “sudden” Type II singularities may or may not be avoided, with the answer being quite sensitive to the polymerization scheme Bojowald (2001); Singh et al. (2006); Cailleteau et al. (2008); Ashtekar (2009); Bamba et al. (2013); Odintsov and Oikonomou (2014, 2016); Chinaglia et al. (2017); Odintsov and Oikonomou (2018); Bojowald et al. (2025). It could also be worth extending these studies to modified gravity models such as $f(R)$ gravity, although the implications for cosmological singularities are likely to depend strongly on the associated effective Hamiltonian and polymerization scheme. While we defer all these points to follow-up work, we stress once more a general but important message of the present work: even ideas or frameworks which at first may appear remote or speculative – such as LQG – can have important consequences for both astrophysical observations and other open problems in physics, whereas the issue of a theoretically consistent formulation can make a significant difference at the level of phenomenological predictions.